OpenStudy (anonymous):

A positive real number is 1 more than its reciprocal. Find the number.

OpenStudy (bahrom7893):

wait sorry it's x = 1/x - 1

OpenStudy (bahrom7893):

x^2 = 1 - x x^2 + x - 1 = 0

OpenStudy (anonymous):

x^2 - x -1 = 0

OpenStudy (bahrom7893):

wait central it's +x, not - x

OpenStudy (anonymous):

no, it's - x x = 1 + (1/x) x^2 = x + 1 x^2 - x -1 = 0

OpenStudy (bahrom7893):

ohh yea, u're right, sorry